OpenStudy (anonymous):

Find the derivative of 6x*arcsin(x)

zepdrix (zepdrix):

\(\Large\bf \color{#CC0033 }{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

Using product rule to set it up, \[\Large\bf\sf \left(6x \arcsin x\right)'\quad=\quad \color{royalblue}{\left(6x\right)'}\arcsin x+6x\color{royalblue}{\left(\arcsin x\right)'}\]

zepdrix (zepdrix):

Just gotta find the derivatives of the blue parts.

zepdrix (zepdrix):

Remember your inverse sine derivative?